The locus of the centre of the circle passing through the intersection of the circles x^2+y^2=1 and x^2+y^2-2x+y=0 is?

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
GIven equations of circles are
x^2+y^2=1 and x^2+y^2-2x+y=0
So by taking a constant\lambdaand writing the equation as
x^2+y^2-1+\lambda(x^2+y^2-2x+y)=0
As given to find out the locus of center of circle we can get its value by substitution of the center
And the locus of intersection is a line with equation as x-2y=0